OpenStudy (anonymous):

The area of a trapezoid is found using the formula , where and are the parallel sides of the trapezoid and h is the height of the trapezoid. What is the area, in square kilometers, of a trapezoid with bases 5 km and 7 km and height 3 km?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Area of a trapezoid = 1/2 * (a+b) * h 'a' and 'b' are the lengths of the parallel sides and 'h' is the height. A = 1/2 * (5 + 7) * (3) = ? square kilometers.